Algebra: How would you plot the second line and why would you plot it like that is there a rule?

X -> 3X+2
-2 = -4
-1 = -1
0 = 2
1 = 5
2 = 8


I get that the first number line will have -2,-1,0,1,2 on it but what about the 2nd one?

Answer by jim_thompson5910(35256)

When x = -2, then 3x+2 = 3(-2)+2 = -6 + 2 = -4

So in short, when x = -2, then 3x+2 = -4

This explains why -2 maps to -4

The same applies to each x value below it.